1.A car is traveling at 90 km/hr. In coming to a stop, it has an acceleration of –8.0 m/s2 . If the reaction
time of the driver is 2.5 s, how far will the car travel before coming to a rest?

2.A speeder passes a parked police car at 35.0 m/s. The police car starts from rest with a uniform
acceleration of 2.60 m/s2
a) How much time passes before the speeder is overtaken by the police car?
b) How far does the speeder get before being overtaken by the police car?

Runner A is initially 4.5 km west of a flagpole and is running with a constant velocity of 5.0 km/h due east.
Runner B is initially 6.5 km east of the flagpole and is running with a constant velocity of 4.5 km/h due west.
What will be the distance of the two runners from the flagpole when their paths cross?
.
1
Expert's answer
2020-10-12T07:49:18-0400

1) "s=\frac{v^2-v_0^2}{2a}=\frac{0^2-25^2}{2\cdot (-8)}=39.1(m)"


2) "s_1=v_0t,s_2=\frac{gt^2}{2}\to s_1=s_2\to v_0t=\frac{gt^2}{2}\to t=2v_0/g=2\cdot35/9.81=7.1(s)"


"s_1=35\cdot7.1\approx250(m)"


3) "x_1+v_1t=x_2+v_2t\to"


"=-4.5+5t=6.5-4.5t\to t=1.158(hours)"


"s=6.5-4.5\cdot1.158=1.29(km)"